An electron starts from rest 41.7 cm from a fixed point charge with -0.133 nC .

How fast will the electron be moving when it is very far away?
Answer #1

chane in potential energy = kinetic energy

kq1q2 / r = 0.5mv^2

9*10^9*0.133*10^-9*1.6*10^-19 / 0.417 = 0.5*9.11*10^-31*v^2

v = 1*10^6 m/s

so the speed of electron will be 1*10^6 m/s